[Brahms-l] Shift report 20050604 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Jun 04 2005 - 07:58:08 EDT
Shifter leader was rk
Shifters were 


1. Had beam from previous shift - due to high background STAR decided to stop taking data and let PHOBOS have collisions. Beam abruptly lost due to power dip at 0358. Recovering since then, injected at 0745.
2. Runs 15170-15176 with FS @ 4deg @ 2442B, MRS @ 60deg @ 700A - 208k tr.2, 287k tr.3.
3. Found DVM off, SMaxR1 off, moved FS to 2.3 deg, changed D2 tap to 1/2 field.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jun 4 07:58:23 2005

This archive was generated by hypermail 2.1.8 : Sat Jun 04 2005 - 07:58:36 EDT